# SteerMusic: Enhanced Musical Consistency for Zero-shot Text-Guided and Personalized Music Editing

### v1

@@ Abstract.
Music editing is an important step in music production, which has broad applications, including game development and film production. Most existing zero-shot text-guided methods rely on pretrained diffusion models by involving forward-backward diffusion processes for editing. However, these methods often struggle to maintain the music content consistency. Additionally, text instructions alone usually fail to accurately describe the desired music. In this paper, we propose two music editing methods that enhance the consistency between the original and edited music by leveraging score distillation. The first method, SteerMusic, is a coarse-grained zero-shot editing approach using delta denoising score. The second method, SteerMusic+, enables fine-grained personalized music editing by manipulating a concept token that represents a user-defined musical style. SteerMusic+ allows for the editing of music into any user-defined musical styles that cannot be achieved by the text instructions alone. Experimental results show that our methods outperform existing approaches in preserving both music content consistency and editing fidelity. User studies further validate that our methods achieve superior music editing quality. Audio examples are available on https://steermusic.pages.dev/.
Text-guided music editing, Personalized music editing, Score distillation, Diffusion model
Work done during the internship at Sony AI
copyright: noneccs: Applied computing Sound and music computingccs: Computing methodologies Neural networks
Figure 1. Steering the music style with text-guided music editing or personalized music editing.
1. Introduction
@@
Text-guided diffusion probabilistic models (DPMs) (Ho et al., 2020; Song et al., [n. d.]) have shown impressive performance in generating diverse and high-quality audio samples, including music, speech, and sound effects. Following the success on language and image domain (Wu et al., 2023; Raffel et al., 2020; Rombach et al., 2022), these text-to-audio diffusion models (Liu et al., 2023a, 2024b; Yang et al., 2023; Deepanway et al., 2023; Majumder et al., 2024) trained with huge datasets can generate diverse samples conditioned on the natural language prompts specified. Consequently, the text-guided music editing task (Wang et al., 2023a; Zhang et al., 2024b) was proposed, which performs music editing by modifying the corresponding text prompts of the text-to-audio generation models, offering greater flexibility in applications. Unlike controllable music generation, music editing task modifies an existing piece of music, which revolves around two primary objectives: preserving original musical content and ensuring alignment between the edited music and the desired target.
Previous methods focus on training music editing models from scratch (Agostinelli et al., 2023; Copet et al., 2023; Wang et al., 2023a) or fine-tuning pretrained text-to-audio models (Paissan et al., 2023; Zhang et al., 2024a), both of which require additional datasets or computational costs. Inspired by recent advances in image editing (Brooks et al., 2023; Meng et al., 2021; Huberman-Spiegelglas et al., 2024; Hertz et al., 2022; Zhang et al., 2021), emerging music editing methods have instead pursued zero-shot techniques to reduce computational overhead. Existing zero-shot text-guided music editing approaches (Zhang et al., 2024b; Manor and Michaeli, 2024; Liu et al., 2024a) introduce noise into the source music during the forward diffusion process to suppress high-frequency components (e.g., timbral information) and subsequently perform editing during the denoising phase based on the target guidance. The process of retrieving a noisy latent representation from the data is commonly referred as inversion step. Due to the imperfect diffusion inversion process, the obtained latent representation may fail to fully preserve the original music content. This issue becomes more serious when the source prompt used as the condition cannot accurately capture the detailed characteristics of the music input (Kawar et al., 2023; Paissan et al., 2023). We refer to the distortion in the inverted latent representation as an “inversion error”. Notably, this distortion occurs even in the reconstruction of a few steps of DDIM inversion (Song et al., [n. d.]) , which changes the original melodic information of the reconstructed results as shown in the CQT spectrogram (Brown, 1991) in Figure 2. In music editing, such distortion can even be compounded, resulting in a failure to preserve instruction-irrelevant content in the original music (Zhang et al., 2024a; Kawar et al., 2023; Meng et al., 2021; Brooks et al., 2023). Although methods such as textual inversion (Gal et al., 2022) have been proposed to mitigate this issue by tuning the textual embeddings for near-lossless reconstruction (Niu et al., 2024), there is no way to manipulate target text prompt within the tuned source textual embeddings in editing tasks. A more promising solution to avoid inversion error is delta denoising score (DDS) (Hertz et al., 2023), a score distillation-based method, which performs editing directly in the data space by computing the difference in denoising scores between the source and target prompts through a single forward step. This approach eliminates the dependency of a full or partial forward diffusion process that could introduce inversion errors. By operating in the data space, DDS enables high-fidelity editing while preserving instruction-irrelevant content on the source data.
@@
Figure 2. How inversion error alters the reconstructed melody in CQT spectrogram after 20 steps DDIM inversion.
@@
Text-guided music editing enables flexible and intuitive modifications, requiring users to provide only an arbitrary text instruction to perform the desired edit. However, one limitation for text-guided music editing is that it often lacks fine-grained control over the direction and nuance of editing. For instance, editing a guitar performance into the one played by person A with a guitar brand B. Text-based editing alone struggles to specify the exact “guitar” required. Moreover, person A and guitar brand B might be unseen concepts for the models, rendering attempts to specify these words in the text prompt ineffective. To address this limitation and enhance user-personalized controllability in text-to-music generation, DreamSound (Plitsis et al., 2024) introduces a pioneering approach that adapts image personalization techniques  (Gal et al., 2022; Ruiz et al., 2023) to the music domain, allowing the extraction of user-defined musical characteristics from reference audio. Besides, DreamSound also demonstrates the potential of leveraging the personalization techniques to perform personalized music editing by manipulating learned musical concepts on the noisy source latent through the denoising process of a personalized diffusion model. Despite their attempts, DreamSound still suffers from the inversion error, and struggles to preserve music content while editing the specific concept given in the text prompt.
@@
  2. (2)
    We propose SteerMusic+, a personalized music editing method that leverages user-defined musical concepts to enable fine-grained customized editing. SteerMusic+, an extension of SteerMusic, enables editing results that are not attainable through text prompts alone. For example, from reggae to the customized reggae given the reference as in Figure 1.
  3. (3)
@@
2. Related work
2.1. Text-guided Music Generation and Editing
In the field of music generation, prior research focuses on approaches that use low-level control signals with strict temporal alignment to the audio outputs. For example, lyrics-conditioned music generation (Yu et al., 2021; Dhariwal et al., 2020) and MIDI-conditioned music generation (Yang et al., 2017). Recently, inspired by advances in natural language processing, high‐level semantic descriptions such as text and visual prompts are gaining popularity in music generation (Dhariwal et al., 2020; Liu et al., 2023a, 2024b; Kundu et al., 2024; Huang et al., 2023a; Chowdhury et al., 2024; Agostinelli et al., 2023; Kreuk et al., 2022; Huang et al., 2023b; Lan et al., 2024; Le Lan et al., [n. d.]; Li et al., 2024; Saito et al., [n. d.]). More recent work also explores melody prompts extracted from reference music (Copet et al., 2024; Novack et al., 2024b, a; Hou et al., 2025; Chen et al., 2024; Kumari et al., 2023), offering more precise and user‐driven music generation capabilities.
Unlike music generation, music editing aims at transforming existing music audio into a target music satisfying the given condition such as text prompts. The goal of music editing is to preserve the original music content while performing the editing. Previous methods involve training music editing models from scratch (Agostinelli et al., 2023; Copet et al., 2023; Wang et al., 2023a; Hou et al., 2025; Mariani et al., 2023) or fine-tuning existing models (Paissan et al., 2023; Zhang et al., 2024a; Tsai et al., 2024; Han et al., 2023; Liu et al., 2023b). Recent methods focus on exploring zero-shot text-guided pipeline (Zhang et al., 2024b; Manor and Michaeli, 2024; Liu et al., 2024a), which rely on pretrained audio generation models. Existing zero-shot text-guided pipelines involve forward-backward diffusion process, which may suffer inversion error as shown in Figure 3 (a).
2.2. Personalized Music Generation and Editing
DreamSound (Plitsis et al., 2024) explores the possibility of capturing musical concepts from the given reference music using a personalization diffusion model (Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023). This method allows users to generate new music samples by incorporating the captured personalized musical concept token into the text prompts. In addition to personalized music generation (Plitsis et al., 2024; Chen et al., 2024), DreamSound further extends their method to personalized music editing. However, we noticed that personalized music editing is still immature. Existing methods (Plitsis et al., 2024) struggle to maintain the musical consistency while editing into the desired music concept captured in the reference audio, which use the same music editing pipeline illustrated in Figure 3 (a).
3. Preliminaries
@@
Figure 3. Overview of two branches of music editing pipeline. (a) is existing music editing pipeline, which involves a inversion process conditioned on source prompt, and perform editing on during denoising; (b) is the proposed pipeline where perform editing on data space by optimizing $x=g(\theta)$. [$S$] indicates a concept token defined by users and gray cycles indicate the optimization trajectory from source to target.
3.1. DPMs on Audio Generation
Given an input audio data $x_{0}$ and the corresponding text prompt $y$, the forward diffusion process is defined as a Markov chain that gradually adds noise to the input audio data $x_{0}$ over $T$ steps, and sampled from marginal distribution $q(x_{t}|x_{0}):=\mathcal{N}(x_{t};\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I)$ as $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$, where $\alpha_{t}$ represents the variance of the forward process at the time step $t$, $\epsilon_{t}\sim\mathcal{N}(0,I)$ at the time step $t$ and $x_{t}$ denotes the noisy latent representation of the input $x_{0}$ at time step $t$. DPMs (Ho et al., 2020) learn the backward diffusion process, which denoises from the prior distribution $\mathcal{N}(0,I)$ to the data distribution $x_{0}$. By predicting the noise $\epsilon_{\phi}(x_{t},y,t)$, the model parameter $\phi$ can be optimized via the following training objective function:
(1) $\mathcal{L}_{\text{DPM}}(x_{0},y;\phi)=\mathbb{E}_{x_{t}\sim q(x_{t}|x_{0}),t\sim\mathcal{U}(1,T)}\|\epsilon_{\phi}(x_{t},y,t)-\epsilon_{t}\|_{2}^{2}$
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is the actual noise added to $x_{0}$, $\mathcal{U}(1,T)$ is a uniform distribution. In audio generation, the diffusion process can be either performed in data space or in a latent space obtained by VAEs (Kingma et al., 2013; Rombach et al., 2022). In this study, we perform music editing using AudioLDM2 (Liu et al., 2024b), a latent diffusion model (LDM) pretrained on Mel-spectrograms. During inference, text prompts are given to AudioLDM2 to generate latent samples. The generated samples are decoded by the pretrained VAE decoder back to the Mel-spectrogram. HiFi-GAN (Kong et al., 2020), a commonly used vocoder, converts the Mel-spectrogram into waveforms.
3.2. Score Distillation and Delta Denoising Score
@@
Score distillation is a technique that leverages the score (i.e., the gradient of the log-density) obtained from a pretrained score-based model (e.g., a DPM $\epsilon_{\phi}$) to refine generated samples, which ensures the alignment of predefined constraints. It is typically implemented via probability density distillation, where gradients from the source diffusion model are used to iteratively refine a differentiable function until the desired outcome is achieved. Score distillation sampling (SDS) (Poole et al., 2022) is the pioneering method in score distillation. Given a differentiable function $x=g(\theta)$ that renders a source input data $x$ with parameters $\theta$, SDS optimizes $\theta$ through a differentiable loss $\mathcal{L}_{\text{Diff}}(\phi,x=g(\theta),y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)-\epsilon\|_{2}^{2}]$. By omitting the UNet Jacobian $\frac{\partial\epsilon_{\phi}}{\partial x_{t}}$ (Poole et al., 2022), the gradient over $\theta$ is given by
@@
(2) $\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\phi,x,y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)-\epsilon)\frac{\partial x}{\partial\theta}]$
where $y^{\text{tgt}}$ is the target prompt, $x_{t}$ is a noisy latent of $x=g(\theta)$ at time step $t$, $w(t)$ is a weighting function. However, SDS is known to suffer from issues such as blurriness in image editing  (Wang et al., 2023b; Hertz et al., 2023). The delta denoising score (DDS) (Hertz et al., 2023) addresses this issue by computing the delta score between the source prompt $y^{\text{src}}$ and the target prompt $y^{\text{tgt}}$. In image editing, DDS refines only the relevant regions of the image corresponding to $y^{\text{tgt}}$, while keeping the other parts of the image unchanged. Given a source input $x^{\text{src}}$ with the corresponding source prompt $y^{\text{src}}$ and the target prompt $y^{\text{tgt}}$, the DDS gradient over the parameters $\theta$ of the differentiable function is
(3) $\nabla_{\theta}\mathcal{L}_{\text{DDS}}(\phi,x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}})=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)\\ -\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t))\frac{\partial x}{\partial\theta}]$
where $x_{t}$ and $x_{t}^{\text{src}}$ shares the same sampled noise $\epsilon$ and time step $t$. See our Supplementary Materials for more detailed background.
@@
While there are different variants of SDS and DDS proposed for image editing (Yu et al., 2025; Hertz et al., 2023; Nam et al., 2024; Lin et al., 2025), the application of DDS in music editing remains underexplored. In the next section, we will explain how to incorporate DDS into music editing.
4. Method
@@
In this section, we introduce two music editing methods. The first method SteerMusic focuses on zero-shot text-guided music editing, and the second method SteerMusic+ focuses on personalized music editing based on a user-defined concept captured from a set of reference music. Unlike the existing forward-backward music editing pipeline in Figure 3 (a), our methods follow a delta denoising framework and perform editing directly on the data space (Figure 3 (b)), which produces results with superior musical content consistency.
4.1. SteerMusic: Zero-shot Text-guided Music Editing
@@
Figure 4. An overview pipeline of SteerMusic+. (a) illustrates a pipeline of fine-tuning a personalized diffusion model given $\mathcal{D}^{\text{ref}}$ and a user-defined [guitar] concept token. (b) provides the detail SteerMusic+ pipeline on personalized music editing that relies on a PDM $\epsilon_{\phi^{\prime}}$ obtained by Eq. 4 in (a). In this example, the PDM captures a user-defined guitar concept. The red dashed lines represent the gradient flows.
@@
We introduce SteerMusic, a zero-shot text-guided music editing method that performs editing in the data space. In this setting, our goal is to edit a source music signal $x^{\text{src}}$ by modifying its corresponding text prompt $y^{\text{src}}$, where $y^{\text{src}}$ is a short textual description of the source music that includes the specific musical attribute intended for modification. The modified text prompt $y^{\text{tgt}}$ acts as the target prompt to guide the editing of $x$. To obtain desirable edited results, the musical content shared by both $y^{\text{tgt}}$ and $y^{\text{src}}$ should be preserved, changing only the contents that are distinct in the $y^{\text{tgt}}$. For example, if the only change in $y^{\text{tgt}}$ compared to $y^{\text{src}}$ is replacing only the word “piano” with “guitar” while keeping the rest of the sentence unchanged, the edited music should preserve the melody and tempo, while modifying only the instrument sound.
@@
To achieve this goal, we adopt the DDS method (Hertz et al., 2023), which has previously been explored only in the image domain. To the best of our knowledge, we are the first to investigate the potential application in music editing. Given a differentiable mapping function $x=g(\theta)$, where we initialize $x=x_{0}^{\text{src}}$, the gradient over $\theta$ is the same as in Eq. 3. In SteerMusic, we set the pretrained diffusion model $\epsilon_{\phi}$ as AudioLDM2 (Liu et al., 2024b). Similar to (Hertz et al., 2023), the delta score in Eq. 3 steers the optimization process toward the target prompt while reducing the noisy editing direction commonly associated with vanilla SDS, leading to enhanced edit fidelity. SteerMusic method enables flexible editing using text instructions alone; however, it is limited to coarse-grained editing, as text prompts often lack the precision information to capture fine-grained musical details.
4.2. SteerMusic+: Personalized Music Editing
@@
As mentioned in Sec. 2, text-guided editing lacks customization for precise music editing, such as transferring music to a specific style. To enable fine-grained personalized editing, we propose SteerMusic+, an extension of SteerMusic. In this setting, we have a set of source music and prompt pair $\{x^{\text{src}},y^{\text{src}}\}$. The target prompt $y^{\text{tgt}}$ is constructed by modifying $y^{\text{src}}$ to manipulate a user-defined concept token $[S]$, representing the desired customization (see example in Figure 4 (b)). SteerMusic+ uses two pretrained diffusion models:
@@
  • A personalized diffusion model (PDM), denoted as $\epsilon_{\phi^{\prime}}$, which is fine-tuned on a small set of dataset contains reference music to capture the user-defined style concept $[S]$ and guide the editing process toward the desired direction.
Inspired by (Yu et al., 2025), we define successful personalized editing based on the following criteria:
@@
In the following parts, we present SteerMusic+ method that enables personalized music editing with enhanced music consistency.

### v3

@@ Abstract
Music editing is an important step in music production, which has broad applications, including game development and film production. Most existing zero-shot text-guided editing methods rely on pretrained diffusion models by involving forward-backward diffusion processes. However, these methods often struggle to preserve the musical content. Additionally, text instructions alone usually fail to accurately describe the desired music. In this paper, we propose two music editing methods that improve the consistency between the original and edited music by leveraging score distillation. The first method, SteerMusic, is a coarse-grained zero-shot editing approach using delta denoising score. The second method, SteerMusic+, enables fine-grained personalized music editing by manipulating a concept token that represents a user-defined musical style. SteerMusic+ allows for the editing of music into user-defined musical styles that cannot be achieved by the text instructions alone. Experimental results show that our methods outperform existing approaches in preserving both music content consistency and editing fidelity. User studies further validate that our methods achieve superior music editing quality.
Codehttps://github.com/sony/steermusic
Demonstration pagehttps://steermusic.pages.dev/
Extended versionhttps://arxiv.org/abs/2504.10826
Introduction
@@
Text-guided diffusion probabilistic models (DPMs) (ho2020denoising; lai2025principles) have shown impressive performance in generating diverse high-quality audio samples, including music, speech, and sounds (mariani2023multi; zhang2024speaking; niu2024soundlocd). These text-to-audio (TTA) diffusion model trained with large scale datasets that can generate diverse samples conditioned on the natural language prompts specified. Consequently, the text-guided music editing task was proposed, which edits music by modifying the corresponding text prompts of the source music. Unlike controllable music generation, music editing task modifies an existing piece of music, which has two primary objectives: preserving original musical content and ensuring alignment between the edited music and the desired target.
@@
Figure 1: The distortion of the reconstructed melody (CQT1-PCC=0.721) after only 20 DDIM inversion steps.
Existing music editing methods focus on training music editing models from scratch (copet2023simple) or fine-tuning pretrained TTA models (zhang2024instruct), both of which require additional datasets or computational costs. Inspired by recent advances in image editing (brooks2023instructpix2pix; huberman2024edit; hertz2022prompt; zhang2021plug), emerging music editing methods have instead pursued zero-shot techniques to reduce computational overhead. Existing zero-shot text-guided music editing pipelines (zhang2024musicmagus; manor2024zero; liu2024medic) introduce noise into the source music during the forward diffusion process to suppress high-frequency components (e.g., timbral information) and subsequently perform editing during the denoising phase based on target guidance in the diffusion latent space. The process of retrieving a noisy latent representation from the data is commonly referred as inversion step. Due to the imperfect diffusion inversion process, the latent representation obtained may not fully preserve the original music content. This issue becomes more serious when the source prompt used as the condition cannot accurately capture the detailed characteristics of the music input (kawar2023imagic; paissan2023audio). We refer to the distortion in the inverted latent representation as an “inversion error”. Notably, this distortion occurs even during the reconstruction of only a few of DDIM inversion steps (song2020denoising) , which alters the melodic information in the reconstructed results compared to the original music as shown in the CQT spectrogram (brown1991calculation) in Fig. 1. In music editing, such distortion can even be compounded, resulting in a failure to preserve instruction-irrelevant content in original music. Although methods such as textual inversion (gal2022image) have been proposed to mitigate this issue by tuning the embeddings for near-lossless audio reconstruction (niu2024soundmorpher), there is no way to manipulate the target text prompts within the tuned source textual embeddings in editing tasks. A more promising solution to avoid inversion error is the delta denoising score (DDS) (hertz2023delta), a score distillation method that performs editing directly in the data space, which defines a differentiable function rendering the source input. DDS computes the difference in denoising scores between the source and target prompts through a single forward step. This approach eliminates the dependency of a full or partial forward diffusion process that could introduce inversion errors. By operating in the data space, DDS enables high-fidelity editing while preserving instruction-irrelevant content on the input.
Figure 2: SteerMusic: Steering the music style with text-guided music editing or personalized music editing.
@@
Text-guided editing enables flexible and intuitive modifications, requiring users to provide only an arbitrary text instruction to perform the desired edit. However, one limitation of text-guided music editing is that it lacks fine-grained control over the direction and nuance of editing. For instance, editing a guitar performance into the one played by person A with a guitar brand B. Text-based editing alone struggles to specify the exact “guitar” required. Moreover, person A and guitar brand B might be unseen concepts for the models, rendering attempts to specify these words in the text prompt ineffective. To enhance user-personalized controllability in text-to-music generation, DreamSound (plitsis2024investigating) introduces a pioneering approach that adapts image personalization techniques (gal2022image; ruiz2023dreambooth) to the music domain, allowing the extraction of user-defined musical characteristics from reference audio. Besides, DreamSound also demonstrates the potential of leveraging the personalization techniques to perform personalized music editing by manipulating learned musical concepts on the noisy source latent through the denoising process of a personalized diffusion model. Despite their attempts, DreamSound still suffers from the inversion error, and struggles to preserve music content while editing the specific concept given in the text prompt.
@@
  2. 2.
    We propose SteerMusic+, a personalized music editing method that leverages user-defined musical concepts to enable customized editing. SteerMusic+, an extension of SteerMusic, enables editing results that are not attainable through text prompts alone. For example, from reggae to the customized reggae given the reference as in Fig. 2.
  3. 3.
@@
Figure 3: Overview of two music editing pipelines: (a) shows the conventional approach, which performs editing during denoising after an inversion process in the diffusion latent space; (b) shows our solution, which directly edits in data space by optimizing the differentiable function $x=g(\theta)$. The differentiable function is initialized with $x^{\text{src}}$. [$S$] denotes a user-defined concept token, and gray circles represent the optimization trajectory from source to target.
Related work
Text-guided Music Generation and Editing
Earlier music generation work focused on low-level control signals with strict temporal alignment, such as lyrics (yu2021conditional; dhariwal2020jukebox) and MIDI (yang2017midinet) conditioning. Recently, high-level semantic prompts have gained popularity (liu2023audioldm; liu2024audioldm; kundu2024emotion; huang2023make; chowdhury2024melfusion; agostinelli2023musiclm; kreuk2022audiogen; huang2023noise2music; le2024high; li2024jen; saitosoundctm). More recent studies further explore melody prompts from reference music (copet2024simple; novack2024ditto; novack2024ditto2; hou2024editing; chen2024jen; kumari2023multi), enabling more precise, user-driven generation.
Music editing transforms existing audio according to target conditions while preserving the original content. Earlier approaches train models from scratch (wang2023audit; agostinelli2023musiclm; copet2023simple; hou2024editing; mariani2023multi) or fine-tune pretrained models (paissan2023audio; zhang2024instruct; tsai2024audio; han2023instructme; liu2023m). Recent work explores zero-shot methods (zhang2024musicmagus; manor2024zero; liu2024medic) using pretrained TTA generator. However, these pipelines rely on a forward-backward diffusion process, which can introduce inversion errors as in Fig. 3 (a).
Personalized Music Generation and Editing
DreamSound (plitsis2024investigating) explores the possibility of capturing musical concepts from the given reference music using a personalization diffusion model (ruiz2023dreambooth; gal2022image; kumari2023multi). This method allows users to generate new music samples by incorporating the captured personalized musical concept token into the text prompts. In addition to personalized music generation (plitsis2024investigating; chen2024jen), DreamSound further extends their method to personalized music editing. However, we noticed that personalized music editing is still immature. Existing methods (plitsis2024investigating) struggle to maintain the musical consistency while editing into the desired music concept captured in the reference audio, which use the same music editing pipeline illustrated in Fig. 3 (a).
Preliminaries
@@
Score distillation refines generated samples using the score (i.e., the gradient of the log-density) from a pretrained diffusion model $\epsilon_{\phi}$ to enforce predefined constraints. It is typically implemented via probability density distillation, where gradients from the source diffusion model are used to iteratively refine a differentiable function until the desired outcome is achieved. Score distillation sampling (SDS) (poole2022dreamfusion) pioneered score distillation by optimizing a differentiable function $x=g(\theta)$ to match a target prompt $y^{\text{tgt}}$, where the function $g(\theta)$ renders the source input $x$ with parameters $\theta$. It minimizes the loss $\mathcal{L}_{\text{Diff}}=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)-\epsilon\|_{2}^{2}]$, where $x_{t}$ is a noisy version of $x$ at time $t$. By omitting the UNet Jacobian, the gradient is $\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\phi,x=g(\theta),y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)-\epsilon)\frac{\partial x}{\partial\theta}]$, where $y^{\text{tgt}}$ is the target prompt, $x_{t}$ is a noisy latent of $x=g(\theta)$ at time step $t$, and $w(t)$ is a weighting function. However, SDS often produces blurry results in image editing  (wang2023prolificdreamer; hertz2023delta). The delta denoising score (DDS) (hertz2023delta) addresses this issue by computing the delta score between the source prompt $y^{\text{src}}$ and the target prompt $y^{\text{tgt}}$.
In image editing, DDS refines only regions relevant to the target prompt $y^{\text{tgt}}$, preserving the rest of the image. Given source input $x^{\text{src}}$ with prompt $y^{\text{src}}$ and target prompt $y^{\text{tgt}}$, the gradient over $\theta$ is
@@
$\nabla_{\theta}\mathcal{L}_{\text{DDS}}(\phi,x=g(\theta),y^{\text{tgt}},x^{\text{src}},y^{\text{src}})\\ =\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t))\frac{\partial x}{\partial\theta}]$ (1)
where $x_{t}$ and $x_{t}^{\text{src}}$ share the same sampled noise $\epsilon$ and the timestep $t$.
@@
Although different variants of SDS and DDS have been proposed for the image domain (yu2025dreamsteerer; hertz2023delta; nam2024contrastive; lin2025zero), the application of DDS in music editing remains underexplored. In the next section, we will explain how to incorporate DDS into music editing tasks.
Method
@@
In this section, we introduce two music editing methods: SteerMusic for zero-shot text-guided editing, and SteerMusic+ for personalized editing using user-defined concepts from reference music. Unlike the forward-backward pipeline in Fig.3 (a), our methods edit directly in a data space (Fig.3 (b)), yielding better musical content consistency.
SteerMusic: Zero-shot Text-guided Music Editing
@@
Figure 4: Overview of the SteerMusic+ pipeline: (a) Personalized diffusion model (PDM) fine-tuned using $\mathcal{D}^{\text{ref}}$ and a user-defined [guitar] concept token. (b) Personalized editing using the PDM $\epsilon_{\phi^{\prime}}$ from (a). Red dashed lines indicate gradient flows.
@@
We introduce SteerMusic, a zero-shot text-guided music editing method that performs editing in the data space. In this setting, our goal is to edit a source music signal $x^{\text{src}}$ by modifying its corresponding text prompt $y^{\text{src}}$, which $y^{\text{src}}$ is a brief description of the source music that includes the specific musical attribute intended for modification. The modified text prompt $y^{\text{tgt}}$ acts as the target prompt to guide the editing. To obtain desirable results, the musical content shared by $y^{\text{tgt}}$ and $y^{\text{src}}$ should be preserved, changing only the content that is distinct in $y^{\text{tgt}}$. For example, if the only change in $y^{\text{tgt}}$ compared to $y^{\text{src}}$ is replacing only the word “piano” with “guitar” while keeping the rest of the sentence unchanged, the edited music should preserve the melody and tempo, while modifying only the musical instrument.
To achieve this goal, we adopt the DDS, which has been previously explored only in the image domain. To the best of our knowledge, we are the first to investigate the potential application in music editing. Following DDS, we define $x=g(\theta)$ rendering a source music signal $x^{\text{src}}$. We picked $g(\theta)=\theta$ as the differentiable function in Eq. 1, where we initialize $\theta=x_{0}^{\text{src}}$. i.e. $x=x_{0}^{\text{src}}$. In SteerMusic, we set $\epsilon_{\phi}$ as a pretrained TTA or text-to-music DPMs.
@@
Similar to hertz2023delta, the delta score in Eq. 1 steers the optimization process toward the target prompt while reducing the noisy editing direction commonly associated with vanilla SDS, leading to enhanced edit fidelity. SteerMusic method enables flexible editing using text instructions alone; however, it is limited to coarse-grained editing, as text prompts often lack the precision information to capture fine-grained musical details.
SteerMusic+: Personalized Music Editing
@@
As mentioned before, text-guided editing lacks customization for precise music editing, such as transferring music to a specific style. To enable fine-grained personalized editing, we propose SteerMusic+, an extension of SteerMusic. In this setting, we have a set of source music and prompt pair $\{x^{\text{src}},y^{\text{src}}\}$. The target prompt $y^{\text{tgt}}$ is constructed by modifying $y^{\text{src}}$ to manipulate a user-defined concept token $[S]$, representing the desired customization (see example in Fig. 4 (b)). SteerMusic+ uses two pretrained diffusion models:
@@
  • A personalized diffusion model (PDM), denoted as $\epsilon_{\phi^{\prime}}$, is fine-tuned on a small set of data containing reference music to capture the user-defined concept $[S]$ and guide the editing process toward the desired direction.
We define successful personalized editing by the criteria:
@@
We now present the SteerMusic+ method that enables personalized music editing with enhanced music consistency.